Given the following confidence interval for a population mean, compute the margin of error, E.

11.81<μ<13.21
Answer #1

Solution :

  = (Lower confidence interval + Upper confidence interval ) / 2

Sample mean = = (11.81+13.21) / 2=12.51

Margin of error = E = Upper confidence interval -   = 13.21 - 12.51=0.7